OpenStudy (anonymous):

A certain isotope decays so that the amount A remaining after t years is given by: A = A0 • e ^(-0.03t) where A0 is the original amount of the isotope. To the nearest year, the half-life of the isotope (the amount of time it takes to decay to half the original amount) is ____ years.

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

the entire thing! I dont even know how to start @Loser66

OpenStudy (loser66):

A = 1/2 A0, solve for t. Dat sit

OpenStudy (loser66):

\(A_0\) is the original amount of the isotope, right?

OpenStudy (loser66):

\( A \) is the new amount of the original one, and it is = a half of original one. That means \(A=\dfrac{1}{2}A_0\)

OpenStudy (loser66):

Now, put them in, you have \(\dfrac{1}{2}A_0=A_0e^{-0.03t}\) simplify \(A_0\) and solve for t. I let you finish the stuff.

OpenStudy (anonymous):

A=A0e^(-0.03t)